Determine the oxidation state of the following:

1. S in Na2SO4
2. H in KOH
3. N in N2O
4. Cu in Cu2S
5. S in H2S2O7
6. Fe in FeS2
7. C in BaCO3
8. F2
9. Ba2+
10.S in SO4^2-
1
Expert's answer
2020-11-24T05:52:29-0500

1. 2+S-8=0 S-6= 0 S= 6

2. 1-2+H= 0 H-1 = 0. H= 1

3. 2N -2 = 0. 2N = 2. N=1

4. 2Cu -2 = 0. 2Cu =2. Cu =1

5. 2+2S-14=0. 2S-12 =0. 2S =12. S=6

6. Fe-4 = 0. Fe =4

7. 2 + C-6 =0 C-4 =0. C=4

8. F2 = 0

9. Ba2+ = +2

10. S -8 =-2. S = -2+8 = 6